OpenStudy (anonymous):

Given the parent functions f(x) = log10 x and g(x) = 3x − 1, what is f(x) • g(x)?

OpenStudy (anonymous):

Please someone help!!

OpenStudy (anonymous):

log(3,(5x - 5)) - log(3,(x - 1)) = log(3,(5x - 5)/(x - 1)) = log(3,5*(x - 1)/(x - 1)) = log(3,5) = ln(5)/ln(3) or about 1.464973520 Given the function f(x) = 0.3(4)x, what is the value of f−1(6)? not sure what f-1(6) means? do you mean inverse as f(x) = 0.3*4 *x = 1.2x y = 1.2x changes to x = 1.2y which means x/1.2 y = x/1.2 as 6/1.2 become 5. Given the parent functions f(x) = 5x − 1 and g(x) = 3x − 9, what is g(x) − f(x)? (3x - 9) - (5x - 1) = 3x - 9 - 5x + 1 = -2x - 8 = -2(x + 4) Given the parent functions f(x) = log10 x and g(x) = 5x − 2, what is f(x) • g(x)? log(10,x) * (5x - 2) as not much you can do excpet maybe log(10,x^(5x - 2)) If f(x) = log2 (x + 4), what is f−1(3)? what is f-1(3)? as I am implying you mean inverse f(x) = log(2,(x + 4)) now inverse is x = log(2,(y + 4)) solve for y x = ln(y + 4) / ln(2) x * ln(2) = ln(y + 4) e^(x * ln(2)) = e^ln(y + 4) e^(ln(2^x)) = e^ln(y + 4) 2^x = y + 4 y = (2^x) - (4) y = (2^3) - (4) y = 8 - 4 y = 4 got it from yahoo answers

OpenStudy (anonymous):

f(x) • g(x) = log10 (3x - 1)x f(x) • g(x) = log10 x3x - 1 f(x) • g(x) = 3x log10 x + log10 x f(x) • g(x) = log10 x - 3x log10 x are the possible answers